- Press Release
- Nov 29, 2022
NASA Space Station On-Orbit Status 4 February 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.
FE Salizhan Sharipov conducted the third experiment session with the Russian/German TEKh-20 Plasma Crystal-3 (PK-3) science payload, modified from yesterday’s run. [First thing in the morning, Salizhan again activated the PK-3/N turbopump in the Service Module (SM) Transfer Compartment (PkhO), which keeps a vacuum inside the work chamber (ZB) in the SM RO (Work Compartment). Then he set up the experiment, supported by tagup with the ground. Video recording began ~19 min after experiment initiation. The turbopump will be deactivated tonight at ~4:25pm EST. The experiment, running in automatic mode, is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside the evacuated work chamber. Main objective today was to obtain highly ordered (crystal) plasma dust at various pressures, particle quantities and disturbances of varying degree.]
CDR/SO Leroy Chiao continued cargo prepacking for the LF-1 Shuttle mission (STS-114), the first RTF (return-to-flight) checkout flight. The activity is supported by an updated prepack list uplinked to the crew. [Return equipment is placed in designated bags and predominantly staged on the FGB floor and in one of the ZSRs (zero-G storage racks) in the Lab (LAB1P4), which is slated to return on LF-1.]
Sharipov had two hours reserved to work on modifying and updating the IMS (Inventory Management System) database to account for transfers to the Progress M-51 (16P) cargo transport, which now doubles as catchall “trash can”. The activity was supported by tagup with ground specialists as required, via S-band.
The FE also conducted an inventory/audit of medical supplies, verifying IMS listings of kits containing psychotropic remedies (PT), preventive remedies (P-2), emergency first aid supplies (NP-2), stowing various medical equipment such as Urolux, Reflotron-4, cardiovascular remedies, anti-inflammatory agents (PV-2), etc., and discarding expired items.
Chiao meanwhile tagged up with MCC-H food specialists via S-band to discuss the results of his recent (2/2) IMS food rations audit.
Salizhan consulted with photo specialists at TsUP/Moscow on the Russian long-term recurring task of imaging PKZ-1V Kromka 1-3 contamination experiment tablet. [The Kromka tablet, deployed on handrail 2614 of the DC-1 “Pirs” docking compartment, collects thruster plume effluents. The pictures were taken in the past with the Kodak 760 digital still camera (DSC) from the EVA hatch 1 “illyuminator” (window) in the DC-1.]
Instead of the regular weekly PCS (Portable Computer System), Chiao today also swapped the HDDs (hard disk drives) of both PCS laptops with spares (#6133, #6036). After their reboot, the ground was to load both PCSs with the new BGA (Beta gimbal assembly) display patch. [The original HDDs (#6026, #6065), which already have the BGA software patch, were stowed as good spares. The intent of this activity was to have two fully configured spare PCS HDDs available for crew use.]
Leroy also rebooted the SSC OCA (Station Support Computer/Orbital Communications Adapter) router laptop, a periodic requirement.
Working off the voluntary Russian task list, the FE conducted the regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment that researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.
The CDR collected the periodic reading of the cabin air’s current CO2 partial pressure in the SM and Lab, using the U.S. CDMK (carbon dioxide monitoring kit, #1009), for calldown (along with the battery status) for use in trending analyses.
Leroy also completed his daily check of the Total Dose reading and End File values of the TEPC (Tissue Equivalent Proportional Counter), which he relocated to the Node on 1/31, and called the data down at the evening DPC (daily planning conference). [This is currently a daily requirement since the UOP (utility outlet panel) near the TEPC’s temporary location (for two weeks) does not allow automated telemetry monitoring by the ground.]
The FE completed the routine SOZh/ECLSS servicing/inspection in the SM, including the ASU toilet facilities. Also included in the maintenance today was the weekly checkup on the BRPK air/liquid condensate separator apparatus. Leroy meanwhile prepared the regular IMS “delta” file for the daily automated export/import to the three IMS databases on the ground.
The crew performed their daily 2.5-hr. physical exercise program on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (both aerobic and anaerobic). Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 1 of a new set) and one hour on VELO.
Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.
At ~2:35pm EST, the crew was scheduled for their regular (nominally weekly) teleconference with the ISS Flight Director at MCC-H.
Energia/TsUP testing of the SM propulsion system (ODU) manifold 1 (both fuel and oxidizer) continued today. [The test is intended to characterize the manifold’s ability to hold a pressure over a period of time. No pressure decay has been observed to date, indicating zero leaks. The test will run for approximately four days, followed by an identical test on manifold 2, scheduled to begin on or about 2/7. This could become a routine activity as determined by the system specialists.]
At ~9:54pm tonight, TsUP/Moscow will remotely conduct a 3-min On/Off test of a reference device (RU), equipped with three laser lights for the docking of the European ATV (Automated Transfer Vehicle) at the SM aft port.
Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were Pyrenees Snowpack, Spain and France (Dynamic Event. This overpass provided an opportunity for photography of snowpack in the Pyrenees mountain range. Looking to the right of track along the northern coastline of Spain; the Pyrenees are the linear east-west trending range that separates the Iberian Peninsula from continental Europe. Heavy snowpack in these mountains will lead to greater runoff and increase the surface water budget for both Spain and France during the year), Phoenix, AZ (this nadir pass provided an opportunity for detailed mapping of the urban core of this rapidly expanding southwestern city. Overlapping mapping swaths of the central urban area are useful for spatial structure mapping, monitoring of infill, and land use change), and Saharan Snow, northern Africa (cold weather conditions in northern Africa have produced snowfall in the Saharan Desert and adjacent mountain ranges. Look to the right of track for the Moroccan Atlas Mountains that parallel the coastline. Images of snow on the peaks and upper slopes are useful for hydrologic studies and assessment of potential geohazards associated with meltwater.
CEO photography can be viewed and studied at the websites:
See also the website “Space Station Challenge” at:
To view the latest photos taken by the expedition 10 crew visit:
- http://spaceflight.nasa.gov/gallery/images/station/crew-10/ndxpage1.html at NASA’s Human Spaceflight website.
Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
Upcoming Key Events:
- Progress M-51 (16P) undocking & destructive reentry — 2/26/05;
- Progress M-52 (17P) launch — 2/28/05.
- EVA-13 — 3/25/05;
- Soyuz TMA-6 (10S) launch — 4/15/05 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
- Soyuz TMA-5 (9S) undock — 4/25/05 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
- Progress M-53 (18P) launch — 6/10/05;
- Progress M-54 (19P) launch — 8/24/05;
- Soyuz TMA-7 (11S) launch — 9/27/05.
ISS Orbit (as of this morning, 7:21am EST [= epoch]):
- Mean altitude — 357.3 km
- Apogee height — 363.8 km
- Perigee height — 350.7 km
- Period — 91.69 min.
- Inclination (to Equator) — 51.64 deg
- Eccentricity — 0.0009738
- Solar Beta Angle — -66.4 deg (magnitude decreasing)
- Orbits per 24-hr. day — 15.70
- Mean altitude loss in last 24 hours — 60 m
- Revolutions since FGB/Zarya launch (Nov. 98) — 35482
ISS Altitude History
Apogee height — Mean Altitude — Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.